OpenStudy (anonymous):

The sum of two-thirds of a number and negative one-ninth is five-sixths. Find the number.

OpenStudy (anonymous):

2/3x -1/9 = 5/6 find x?

OpenStudy (sohailiftikhar):

\[2/3x-1/9x=5/6\]

OpenStudy (sohailiftikhar):

now can you simplify the terms containing x ?

OpenStudy (anonymous):

i think so!

OpenStudy (sohailiftikhar):

\[(3(2)x-1x)\div9\]

OpenStudy (madhu.mukherjee.946):

x=51/36

OpenStudy (sohailiftikhar):

don't tell the answer please ...directly

OpenStudy (madhu.mukherjee.946):

2x/3-1/9=5/6 6x-1/9=5/6

OpenStudy (anonymous):

okay the answer sheet says 2 5/6? so i have to turn it into a mixed fraction>

OpenStudy (madhu.mukherjee.946):

36x-6=45

OpenStudy (madhu.mukherjee.946):

why would 1/9 be multiplied with x

OpenStudy (sohailiftikhar):

oh ! yes you are right

OpenStudy (madhu.mukherjee.946):

so...............

OpenStudy (anonymous):

where did the 6x come from i am lost

OpenStudy (madhu.mukherjee.946):

dude again lcm

OpenStudy (anonymous):

Oh yeah okay

OpenStudy (madhu.mukherjee.946):

wait i'm showing you

OpenStudy (madhu.mukherjee.946):

you find the lcm of 3and 9 its 9

OpenStudy (madhu.mukherjee.946):

so we are multiplying 2x with 3

OpenStudy (madhu.mukherjee.946):

you get that?

OpenStudy (anonymous):

yeah i get it lol

OpenStudy (anonymous):

thanks !

OpenStudy (sohailiftikhar):

your welcome

OpenStudy (madhu.mukherjee.946):

whts this

OpenStudy (sohailiftikhar):

6x-1/9=5/6 6x-1=(5/6)*9 6x-1=45/6 6x-1=15/2 6x=15/2+1 6x=15+2/2 6x=17/2 x=17/12

OpenStudy (madhu.mukherjee.946):

6x-1=15/2 6x=15/2+1 6x=17/2 x=17/12

OpenStudy (sohailiftikhar):

yes

OpenStudy (sohailiftikhar):

:)

OpenStudy (madhu.mukherjee.946):

okay i got confused:p sorry

OpenStudy (madhu.mukherjee.946):

actually 51/36=17/12

OpenStudy (sohailiftikhar):

lol 51/36=17/12 you just missed 3 with 6 you wrote 51/6 it's ok :)

OpenStudy (madhu.mukherjee.946):

so i'm correct

OpenStudy (anonymous):

thanks guys

OpenStudy (sohailiftikhar):

yes you are :) and np @glitterythings